Four bar chain is given with length of links as follow:


Link 1 = 74mm, link 2 = 96 mm, link 3 = 60 mm, Link 4 = 70. Determine:


a) Whether it belong to class 1 or class 2 category of mechanisms.


b) Describe all possible inversions with explanations.

1
Expert's answer
2022-02-24T18:40:02-0500

Link 1=74mm=b

Link 2=96mm=a

Link 3=60mm=d

Link 4=70mm=c

for class 1, (a-b)<(c-d)

for class 2, (a-b)>(c-d)

for this system: (a-b)=(96-74)=22

(c-d)=(70-60)=10

hence, (a-b)>(c-d) thus it is a class 2 category of mechanism.